OpenStudy (anonymous):

Write the expression below as a single real number. Do not use decimal approximations. sin^2(-π/3)cos(7π/3)-2tan(-π/6)

OpenStudy (anonymous):

I wrote: \[(-\sqrt{3}/2)^2(1/2)-2(-1/\sqrt{3})\]

OpenStudy (jamesj):

Looks good, just simplify that expression a little.

OpenStudy (anonymous):

\[Answer = \frac{16\sqrt{3}+9}{24}\]

OpenStudy (anonymous):

Just looked weird and I wanted to make sure I was doing my math right.